OpenStudy (anonymous):

please help me :) !!!!! Find f(6) if f(x) =1/2 x + x - 4

OpenStudy (16shuston):

plug 6 into x

OpenStudy (anonymous):

so it would be 6x 6?

OpenStudy (16shuston):

so it would be 1/2(6)+6-4 solve

OpenStudy (anonymous):

okay so 1/2 x 6= 7.2 and 6+4=10

OpenStudy (16shuston):

no, 1/2*6=3+6-4=?

OpenStudy (anonymous):

the answer i got is 5 ?

OpenStudy (16shuston):

yes u r correct

OpenStudy (16shuston):

do u understand the process?

OpenStudy (anonymous):

really thats the right answer :)

OpenStudy (anonymous):

i kinda get it.

OpenStudy (16shuston):

ya u r 100% correct.. okay all u do is whenever there is an x u replace it with the number u r looking for like for example f(10) whenever there is an x replace it with 10

OpenStudy (anonymous):

wow i didnt think of it like that :) that really helps! hey thanks for your help! God Bless You:)

OpenStudy (16shuston):

your welcome, im here anytime except for after this time of night because im going to bed. but other than that, fan me and you will know when im on and i would love to help u!!

OpenStudy (anonymous):

thanks!:)